OpenStudy (thomas5267):

Solve \(2^{0.5x}=3^{-0.5x}+\frac{5}{3}\).

OpenStudy (thomas5267):

@Kainui The answer is 2 but I couldn't solve it without using graphic display calculator.

OpenStudy (thomas5267):

@ganeshie8 @ikram002p @Marki

OpenStudy (anonymous):

yes its 2

OpenStudy (thomas5267):

But how to solve it without using calculator?

OpenStudy (anonymous):

do u have to give an explanation

OpenStudy (thomas5267):

It was a past paper question in a paper which prohibits the use of calculator.

OpenStudy (anonymous):

i dont know then really im sorry

OpenStudy (mathmath333):

\(\large \color{black}{ 2^{0.5x}=3^{-0.5x}+\frac{5}{3} \hspace{.33em}\\~\\ 2^{x/2}=3^{-x/2}+\frac{5}{3} \hspace{.33em}\\~\\ 2^{x/2}-3^{-x/2}=\frac{5}{3} \hspace{.33em}\\~\\ 2^{x/2}-\dfrac{1}{3^{x/2}}=\frac{6-1}{3} \hspace{.33em}\\~\\ 2^{x/2}-\dfrac{1}{3^{x/2}}=2-\frac{1}{3} \hspace{.33em}\\~\\ \normalsize \text{from this point it is evident that}\quad x/2=1\implies x=2\hspace{.33em}\\~\\ }\)

ganeshie8 (ganeshie8):

Neat :)

OpenStudy (thomas5267):

So there is no general method to solve this kind of equations? For exmaple, does \(2^{0.5x}=3^{-0.5x}+7\) have a nice solution?